POSSIBLE OBSERVATION OF THE ^{3}A′ - ^{1}A′ ELECTRONIC TRANSITION OF THE METHYLENE PER-OXY CRIEGEE INTERMEDIATE

\underline{NEAL D. KLINE} and TERRY A. MILLER, Laser Spectroscopy Facility, Department of Chemistry, The Ohio State University, 120 W. 18th Avenue, Columbus, Ohio 43210; MARC COONS and JOHN HERBERT, Department of Chemistry, The Ohio State University, Columbus, Ohio 43210.

It is possible that we have observed for the first time the transition from the ground ^{1}A′ state to the low lying ^{3}A′ state of the methylene peroxy Criegee intermediate, CH_{2}O_{2}, which is otherwise known as carbonyl or formaldehyde oxide. Criegee intermediates are reactive intermediates that are formed in the ozonolysis of olefins in both liquid and gaseous phases with the methylene peroxy intermediate being formed specifically from the ozonolysis of ethene. In the atmosphere Criegee intermediates are formed in reactions that lead to secondary organic aerosols and participate in reactions with SO_{2} and NO_{2}. A peroxy-like spectrum was observed in the near-IR using cavity ringdown spectroscopy after photolysis of a diiodomethane precursor at 248 nm followed by a reaction with O_{2}. Possible assignment of the spectrum to CH_{2}O_{2} is based on a strong analogy between the electronic structure of methylene peroxy and ozone. However experiments and electronic structure calculations are continuing to positively attribute the spectrum to either CH_{2}O_{2} or CH_{2}IO_{2} which could also be formed by this chemistry.